# Classical optimization with imaginary time block encoding on quantum computers: The MaxCut problem

### v1

@@ III.2 QAOA + ITE-BE
Figure 4: The averaged approximation ratio $r$, probability of obtaining optimal solution $p_{\rm opt}$, and the post-selection success rate within three standard deviations of the mean for the QAOA+ITE-BE method. Here we use $p-$level QAOA circuit for state preparation, where $p=4,6,8,10,12$. The QAOA input state leads to faster convergence of ITE-BE and a post-selection rate that is one order of magnitude higher than using an equal superposition state to solve same graph for $p=12$.
Our second method combines the QAOA with the imaginary time evolution (see Fig. 2). In the first part of the protocol, instead of training QAOA for parameter $(\vec{\beta},\vec{\gamma})$, we adopt near-optimal values from Ref. [29] and prepare the input state using a $p-$level QAOA circuit. The QAOA output state has a higher overlap with the ground state than the standard equal superposition state $|+\rangle^{\otimes n}$. As such we expect this protocol to improve the post-selection success rate in the subsequent ITE-BE part. In the second part, we apply the ITE-BE method to further search for the optimal solution, where each imaginary time propagator for edge $\langle j,k\rangle$ is implemented by the circuit block in Fig. 1(a). To reduce the resource overhead of the ITE-BE part, we reuse ancilla qubits for each circuit block by mid-circuit measurement and reset as in the pure ITE-BE method. The data qubits are measured only after all circuit blocks have been executed.
In order to evaluate its performance, we apply this hybrid approach to solve the MaxCut problem on the same set of randomly generated unweight 3-regular(u3R) graph instances as in Sec. III.1, following the same settings. We report the $99.7\%$ averaged uncertainty interval for the averaged approximation ratio $r$, probability of optimal $p_{\rm opt}$ and post-selection success rate as a function of imaginary time $\tau$ for graphs with $N=12$ and different QAOA levels in Fig. 4. The black dashed line is the classical bound of approximation ratio $r\approx 0.9326$.
@@ IV Conclusion and Outlook
In this paper, we propose two quantum algorithms based on the block-encoding implementation of the imaginary time evolution propagator (ITE-BE) and evaluate their performance on the MaxCut problem. We note that ITE-BE requires no parameter optimization and can be applied to any Pauli string Hamiltonian, see Fig. 1 for illustrations. We conduct numerical simulations on randomly generated unweight $3-$regular graphs with vertex number from $N=6$ to $N=12$. Our result shows that both ITE-BE based methods guarantee convergence to the ground state when the imaginary time $\tau$ is sufficiently large, and the ITE-BE based methods are promising to solve combinatorial optimization problems with high accuracy at the cost of larger number of shots due to post-selection requirements.
The success of QAOA+ITE-BE method emphasizes the importance of concatenating quantum algorithms in order to boost overall performance and overcome individual limitations. When preparing the initial state with a $p-$level QAOA for the ITE-BE layer, our results show a faster convergence of ITE-BE to the optimal solution and a post-selection rate that is an order of magnitude higher than using $|+\rangle^{\otimes n}$ as initial state. The result of QAOA+ITE-BE also reveals a trade-off between the QAOA level $p$ and the imaginary time $\tau$, namely, one can choose a lower QAOA level with a large $\tau$ to ensure convergence using a shallow circuit, or choose higher QAOA level to have quicker convergence with higher post-selection rate. Moreover, integrating QAOA with the ITE-BE method also benefits QAOA, as it avoids the need for additional training of QAOA with near-optimal parameters to achieve better accuracy. Instead, the subsequent ITE-BE layer enhances the result with only $O(E)$ circuit depth, where $E$ is the number of edges in the graph. This improvement is particularly significant when solving MaxCut on large graphs.
@@ V Acknowledgment
A.F. acknowledges funding by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research Quantum Testbed Program under contract DE-AC02-05CH11231. The authors acknowledge the Center for Advanced Research Computing (CARC) at the University of Southern California for providing computing resources that have contributed to the research results reported within this publication. URL: https://carc.usc.edu.
@@ Appendix B The MaxCut Solution of Bipartite Graphs
Figure 6: The averaged performance of QAOA+ITE-BE method on 40 randomly generated bipartite $3-$regular graphs for $N=6,8,10$ and $12$. In this case, the QAOA initial state is very close to the optimal solution so the ITE-BE method have only smaller improvement. The success rate reaches a plateau after imaginary time $\tau\sim 0.4$.
During the process of data analysis, we observe that both ITE-BE methods converge more quickly for some graphs than for others in our sample. Moreover, these graphs have higher post-selection success rate compared to other graph with same size, and the success rate reaches a plateau after a certain imaginary time $\tau$. To understand the conditions under which the ITE-BE method converges more quickly, we compute the spectra of both the adjacency matrices and Hamiltonians of these special graphs. We found that the minimum eigenvalue of the adjacency matrices of these graphs is $-3$, which is a characteristic feature of unweighted bipartite $3-$regular graphs. In addition, the minimum eigenvalue of their Hamiltonian $H_{C}$ is given by $(\frac{3N}{2}-1)/2$, where $\frac{3N}{2}$ is the total number of edges in a $3-$regular graph. This implies that the MaxCut solution is a partition where all edges in the graph connect the two parts of the partition. These observations conclusively identify the special graphs as bipartite $3-$regular graphs.
@@
It is worth noting that for any bipartite graph with weights $w_{jk}=1$, the solution to the MaxCut problem is unique and corresponds to the bipartition of the graph. This is because, in a bipartite graph, all edges connect vertices from one part to the other, so the total weight of the edges crossing the bipartition is maximized and cannot be increased further. In this case, the MaxCut solution can be found using classical algorithms such as breadth-first search (BFS) or depth-first search (DFS) with linear time complexity $O(V+E)$. Therefore, a quantum algorithm cannot achieve exponential speed-up to provide a quantum advantage for MaxCut problem in bipartite graph, but should not perform worse than classical algorithms.
Figure 7: The averaged performance of QAOA+ITE-BE method on 40 randomly generated bipartite $3-$regular graphs for $N=6,8,10$. The ITE-BE layer further enhances the performance of QAOA especially for large graphs.
@@ Appendix C QAOA+ITE-BE Results for Graphs with 6, 8, 10 Nodes
Appendix D Approximation Ratio
Here is an illustration of why the approximation ratio for $|+\rangle^{\otimes n}$ is above $0.5$. For an arbitrary state $|\psi\rangle$, the cost function of $\langle H_{C}\rangle$ where $w_{jk}=1$ is given by
$\langle\psi|H_{C}|\psi\rangle=\frac{1}{2}\sum_{\langle j,k\rangle}\left(\langle\psi|Z_{j}Z_{k}|\psi\rangle-\langle\psi|I|\psi\rangle\right).$ (14)
There are totally $23N/$ 2 edges for a $3-$regular graph, so
$\langle H_{C}\rangle=\frac{1}{2}\sum_{\langle j,k\rangle}\langle\psi|Z_{j}Z_{k}|\psi\rangle-\frac{3N}{4}.$ (15)
If consider bipartite graph and let $|\phi\rangle=\frac{1}{\sqrt{2^{n}}}|+\rangle^{\otimes n}$, then the first term $\frac{1}{2}\sum_{\langle j,k\rangle}\langle\psi|Z_{j}Z_{k}|\psi\rangle$ is very close to 00, so $\langle H_{C}\rangle_{\phi}=\frac{3N}{4}$. The MaxCut for bipartite graph will cut all edges of the graph, so $\langle\psi|Z_{j}Z_{k}|\psi\rangle=-1$ for any $j,k$. Then for optimal solution,
$\langle H_{C}\rangle=\frac{1}{2}\sum_{\langle j,k\rangle}\left(\langle\psi|Z_{j}Z_{k}|\psi\rangle\right)-\frac{3N}{4}=-\frac{3N}{4}-\frac{3N}{4}=-\frac{3N}{2},$ (16)
and the approximation ratio for $|\phi\rangle$ on bipartite graph is
$r=\frac{\langle H_{C}\rangle_{\phi}}{\langle H_{C}\rangle}=0.5.$ (17)
For non-bipartite graph, more than one $\langle\psi|Z_{j}Z_{k}|\psi\rangle=1$ in the sum of eq. 15. In this case, the cost function for the MaxCut will still be negative but larger than $-\frac{3N}{2}$, then the approximation ratio for $|\phi\rangle$ on non-bipartite graph is
$r^{\prime}=\frac{\langle H_{C}\rangle_{\phi}}{\langle H_{C}\rangle^{\prime}}\geq\frac{\langle H_{C}\rangle_{\phi}}{\langle H_{C}\rangle}=0.5.$ (18)

### v2

@@ III.2 QAOA + ITE-BE
Figure 4: The averaged approximation ratio $r$, probability of obtaining optimal solution $p_{\rm opt}$, and the post-selection success rate within three standard deviations of the mean for the QAOA+ITE-BE method. Here we use $p$-level QAOA circuit for state preparation, where $p=4,6,8,10,12$. The QAOA input state leads to faster convergence of ITE-BE and a post-selection rate that is one order of magnitude higher than using an equal superposition state to solve same graph for $p=12$. As comparison, we show the results from purely ITE-BE method labeled as $p=0$ in above figures, which indicates that the ITE-BE has faster convergence to the optimal solution and nearly $10$ times higher post-selection rate when using $p$-level QAOA input state.
Our second method combines the QAOA with the imaginary time evolution (see Fig. 2). In the first part of the protocol, we directly use near-optimal values $(\vec{\beta},\vec{\gamma})$ presented in Ref. [49] without any further training of QAOA to prepare an input state for the ITE-BE part. The QAOA output state has a higher overlap with the ground state than the standard equal superposition state $|+\rangle^{\otimes n}$. As such we expect this protocol to improve the post-selection success rate in the subsequent ITE-BE part. In the second part, we apply the ITE-BE method to further search for the optimal solution, where each imaginary time propagator for edge $\langle j,k\rangle$ is implemented by the circuit block in Fig. 1(a). To implement the overall protocol for a given undirected graph $G=(V,E)$, we could reuse ancilla qubits for each circuit block in the ITE-BE part such that only $|V|+1$ qubits is required. The data qubits are measured only after all circuit blocks have been executed. Also, the circuit depth of $e^{-i\gamma_{k}H_{C}}$ in state preparation is $O(|E|)$, so the circuit for overall protocol has $O(p|E|+|E|)$ depth.
In order to evaluate its performance, we apply this hybrid approach to solve the MaxCut problem on the same set of randomly generated unweighted 3-regular(u3R) graph instances as in Sec. III.1, following the same settings. We report the $99.7\%$ averaged uncertainty interval for the averaged approximation ratio $r$, probability of optimal $p_{\rm opt}$ and post-selection success rate as a function of imaginary time $\tau$ for graphs with $N=12$ and different QAOA levels in Fig. 4.
@@ IV Conclusion and Outlook
In this paper, we propose two quantum algorithms based on the block-encoding implementation of the imaginary time evolution propagator (ITE-BE) and evaluate their performance on the MaxCut problem. We conduct numerical simulations on randomly generated unweighted 33-regular graphs with vertex number from $N=6$ to $N=12$. Using circuits with $N$ qubits and $O(|E|)$ depth, our results show that both ITE-BE based methods converge to the ground state when the imaginary time $\tau$ is sufficiently large. This performance surpasses that of the QAOA algorithm. In comparison to classical MaxCut solvers, current semi-definite programming approaches can solve dense graph instances (density up to $0.9$) for up to $N=100$ nodes (BiqMac [54]) or $N=180$ nodes (BiqBin [55]). Moreover, modern linear programming solver can handle sparse graphs with hundreds or thousands of nodes [56], and even more than N=10, [^0] N=10{,}000 nodes (McSparse [57]). Thus, demonstrating quantum advantage would require access to more than $100$ qubits. However, the guaranteed convergence of the ITE-BE based methods still indicates their promising potential for solving MaxCut problems on large graphs.
The success of QAOA+ITE-BE method emphasizes the importance of concatenating quantum algorithms in order to boost overall performance and overcome individual limitations. When preparing the initial state with a $p$-level QAOA for the ITE-BE layer, our results show a faster convergence of ITE-BE to the optimal solution and a post-selection rate that is an order of magnitude higher than using $|+\rangle^{\otimes n}$ as initial state. The result of QAOA+ITE-BE also reveals a trade-off between the QAOA level $p$ and the imaginary time $\tau$, namely, one can choose a lower QAOA level with a large $\tau$ to ensure convergence using a shallow circuit, or choose higher QAOA level to have quicker convergence with higher post-selection rate. Moreover, integrating QAOA with the ITE-BE method also benefits QAOA, as it avoids the need for additional training of QAOA with near-optimal parameters to achieve better accuracy. Instead, the subsequent ITE-BE layer enhances the result with only $O(|E|)$ circuit depth, where $|E|$ is the number of edges in the graph. This improvement is particularly significant when solving MaxCut on large graphs.
@@ V Acknowledgment
This research was supported by the U.S. Department of Energy (DOE) under Contract No. DE-AC02-05CH11231, through the National Energy Research Scientific Computing Center (NERSC), an Office of Science User Facility located at Lawrence Berkeley National Laboratory. A.F. acknowledges funding by the U.S. Department of Energy, Office of Science, Office of Advanced Scientific Computing Research Quantum Testbed Program under contract DE-AC02-05CH11231. The authors acknowledge the Center for Advanced Research Computing (CARC) at the University of Southern California for providing computing resources that have contributed to the research results reported within this publication. URL: https://carc.usc.edu.
@@ Appendix B The MaxCut Solution of Bipartite Graphs
Figure 6: The averaged performance of QAOA+ITE-BE method on 40 randomly generated bipartite $3$-regular graphs for $N=6,8,10$ and $12$. In this case, the QAOA initial state is very close to the optimal solution so the ITE-BE method have only smaller improvement. The success rate reaches a plateau after imaginary time $\tau\sim 0.4$. The $p=0$ curves give the results from purely ITE-BE method on same bipartite graphs.
During the process of data analysis, we observe that both ITE-BE methods converge more quickly for some graphs than for others in our sample. Moreover, these graphs have higher post-selection success rate compared to other graph with same size, and the success rate reaches a plateau after a certain imaginary time $\tau$. To understand the conditions under which the ITE-BE method converges more quickly, we compute the spectra of both the adjacency matrices and Hamiltonians of these special graphs. We found that the minimum eigenvalue of the adjacency matrices of these graphs is 3-3, which is a characteristic feature of unweighted bipartite $3$-regular graphs. In addition, the minimum eigenvalue of their Hamiltonian $H_{C}$ is given by $(\frac{3N}{2}-1)/2$, where $\frac{3N}{2}$ is the total number of edges in a 33-regular graph. This implies that the MaxCut solution is a partition where all edges in the graph connect the two parts of the partition. These observations conclusively identify the special graphs as bipartite $3$-regular graphs.
@@
It is worth noting that any bipartite graphs with non-negative weights and only one connected component, the solution to the MaxCut problem is unique up to a “flipping” of all vertices and corresponds to the bipartition of the graph. This is because, in a bipartite graph, all edges connect vertices from one part to the other, so the total weight of the edges crossing the bipartition is maximized and cannot be increased further. In this case, the MaxCut solution can be found using classical algorithms such as breadth-first search (BFS) or depth-first search (DFS) with linear time complexity $O(V+E)$.
Figure 7: The averaged performance of QAOA+ITE-BE method on 100 randomly generated u3R graphs for $N=6,8,10$. The ITE-BE layer further enhances the performance of QAOA especially for large graphs. We also show results obtained from purely ITE-BE method ($p=0$) on same u3R graph instances as a comparison.